OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is 3/4 cup, how many servings are in the box?

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

on serving is less than a cup, so for one thing there must be more than ten servings in a ten cup box!

OpenStudy (anonymous):

correct answer is \[10\div \frac{3}{4}=10\times \frac{4}{3}=\frac{40}{3}=13\tfrac{1}{3}\]

OpenStudy (anonymous):

as usual here, caveat emptor